Help PLEASE!
3. Find M.M.E for the parameter 0 There are 3 observations, X1 3(1-0) = 1,X2 2,X3 = 1 20 P(X1 1) = P(X1 2) = 3 - e 3-0
Answer #1

The solution for MME of \theta is as follows:

net X be a r-re. representing the dirth of X, X2 ,XBrie, X, X2 X3 id fx (DC) ...P[x=d] = 301-0) P[x=2] - 200 E(x) = I x. P [x